Mike wants to put R1, 000,000 investment/capital in the above project. He has analysed the estimated income and expenditure and found that the project will give incremental net cash flows or net operational cash flows of R360, 000 per year. The project has a lifespan of 5 years (salvage value is ignored). The expected average annual depreciation is R4000. Use the discount rate you calculated above to determine the Pay Back Period, Rate of Return, NPV and the IRR of the project and comment about your results.

- Pay Back Period

- Return On Investment

- Net Present Value

- IRR


1
Expert's answer
2021-10-14T13:15:12-0400

Find the discount rate using the formula:

Rd=d+rmre+rliq+rtec+rman

d - risk-free rate

rmre - compensation for industry risk, for example, the risk of investing in the real estate market;

rliq - compensation for liquidity risk;

rtec - compensation for technogenic risks;

rman - compensation for the risk of poor management

let then

d=5%

rmre=2%

rliq=3%

rtec=1%

rman=1%

Rd=5+2+3+1+1=12%

We present the calculated data:


"NPV=-I+\frac{CF}{(1+rd)^1}+...\frac{CF}{(1+rd)^n}=-1000 000+\frac{360 000}{(1+0,12)^1}+\frac{360 000}{(1+0,12)^2}+\frac{360 000}{(1+0,12)^3}+\frac{360 000}{(1+0,12)^4}+\frac{360 000}{(1+0,12)^5}=297719.43"


"NPV=-I+\frac{CF}{(1+irr)^1}+...\frac{CF}{(1+irr)^n}=0"


"297719.43=-I+\frac{360 000}{(1+irr)^1}+...\frac{360 000}{(1+irr)^5}=0"

IRR=10%


"payback period =4+\frac{204 273.67}{228786.51}=4.89"

The payback period is the time it takes to recover the value of the initial investment, that is, the time when the investment reaches the break-even point. It calculates the number of years it takes us to get the initial investment value. Therefore, we take the number of years when we received a positive cash flow and add the ratio of the cash flow of the last year to the previous one.


We consider ROI as the ratio of net cash flow to investment


"ROI=\frac{360\times5}{1 000 000}=1.8"




Net Present Value> 0, the project can be run